OpenStudy (anonymous):

In a bin containing 500 marbles, 160 are red, 220 are blue, and 120 are green. In a bowl of 25 marbles scooped from this bin, how many would be expected to be blue? A. 6 B. 8 C. 11 D. 22

rebeccaxhawaii (rebeccaxhawaii):

add up all the marbles - whats 500+260+220+120=?

OpenStudy (anonymous):

ik this is easy im in a hurry tho

rebeccaxhawaii (rebeccaxhawaii):

well we are here to help, not give out answers

OpenStudy (abdullahm):

There are 500 total marbles. The question states this. 220 out of the 500 are blue. We want to know how many will be blue out of 25. We need to make a proportion: x/25 = 220/500 Do you know how to cross multiply?

OpenStudy (anonymous):

thats not what im askin for.sorry if you miss understood that

OpenStudy (anonymous):

d?

OpenStudy (debbieg):

Does it seem reasonable that 22 out of 25 would be blue, if only 220 out of 500 are blue? That's less than half of the total marbles that are blue.

OpenStudy (abdullahm):

cross multiply it \(\Large\bf \frac{a}{b} =\frac{c}{d}\) becomes \(\large\bf ad = bc\)